hi
we know that ae≠bd and I want to find the solutions, I mean(x,y), of this equations
ax2+by3=c
dx2+ey3=f
I solved it by Matrix, and I think that this system of equation has three answers, I will be very grateful to you if you solve and explain this to me.
thanks.

Answer by ikleyn(52858)

.
Introduce new variables u = x^2 and v = y^3.

Solve the linear system for these new variables (using determinants).

Then return to the original variables x and y.


See the lessons
    - Solving systems of non-linear equations in two unknowns using the Cramer's rule
    - Solving systems of non-linear equations in three unknowns using Cramer's rule
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of equations that are not linear".